# RePO: ReLU-based Preference Optimization

### v1

@@ 3.4 Over-Optimization Analysis
When a larger $\gamma$ is selected, RePO permits most data to bypass the filter, resulting in a continuous increase in the reward margin between winning and losing samples. This phenomenon, a manifestation of reward exploitation or hacking (Amodei et al., 2016), is often referred to as “over-optimization” in the RLHF literature, and has been empirically investigated in both controlled experiments (Gao et al., 2023) and user studies (Dubois et al., 2023).
Model Over-Optimization: Building on Rafailov et al. (2024), we investigate over-optimization in RePO, by evaluating six different values of $\gamma$ ($0.0,0.2,0.4,0.6,0.8,1.0$), each corresponding to varying levels of data filtering. Across all cases, we observe a distinct hump-shaped performance pattern: while moderate filtering improves alignment, excessive filtering causes performance to degrade, highlighting the over-optimization effect.
Figure 6: Over-optimization results for RePO on Llama3-8B-IT and Gemma2-9B-IT. Metrics include model LC win rate on AlpacaEval2 and raw win rate on Arena-Hard. Fitted dotted curves apply scaling laws from Gao et al. (2023) to direct alignment, with GPT-4 win rates substituting gold reward model scores.
Scaling Law Fits. Previous work (Gao et al., 2023; Rafailov et al., 2024) has established scaling laws for reward model scores as a function of the KL divergence between the initial and optimized policies. In contrast, we eliminate the reference model and the associated computational cost of calculating KL divergence. Instead, we use the mean implicit reward margin during training as a proxy metric. The reward function $R(d)$ is given by:
@@
3.5 RePO++
Building on the previous discussion, we identify the advantages of RePO in finding a “cutoff point” to control over-optimization. However, a limitation persists: for cases where the implicit reward margin is smaller than $\gamma$, their gradient weights become unit, failing to differentiate between samples.
One potential solution is to apply a non-linear activation function to weight the gradient of each pair. Specifically, we extend RePO to DPO and SimPO, introducing a new method, RePO++.
@@ 3.5 RePO++
What does the RePO++ update do? We again consider analyzing the gradient of the loss function RePO++ for an intuitive understanding. The gradient w.r.t. the parameters $\theta$ is given by:
@@
where $s_{\theta}=\sigma\left(\beta\left(-M_{\theta}+\gamma\right)\right)$ and $\mathbb{I}_{\theta}$ is an indicator function that is 1 if $M_{\theta}<\gamma$ and 0 otherwise. Intuitively, This gradient combines the strengths of SimPO and RePO: it scales updates by $s_{\theta}$ and filters them using $\mathbb{I}_{\theta}$, ensuring that the model focuses on less-separated pairs, with smaller separations receiving higher weights.
@@
4 Theoretical Analysis of RePO
Next, we relate preference optimization to a supervised binary classification task and show how RePO leverages the convex envelope of the 0-1 loss to achieve both theoretical optimality and efficient optimization.
Following Tang et al. (2024), we reformulate reward modeling as a supervised binary classification task. Given data pairs $(z,l)$ with $z\in\mathbb{R}^{k}$ and labels $l\in\{-1,1\}$, the goal is to learn a predictor ^$\hat{\ell}(z)\in\mathbb{R}$ such that $\mathrm{sign}(\hat{\ell}(z))$ matches $l$. The classification accuracy is given by: $\frac{1}{2}\mathbb{E}[\mathrm{sign}(\hat{\ell}(z)\cdot l)]+\frac{1}{2},$ and the corresponding loss function is:
$\mathcal{L}_{0\text{-}1}(\hat{\ell}):=\mathbb{E}\left[1-\mathrm{sign}(\hat{\ell}(z)\cdot l)\right].$
The above loss, known as the 0-1 loss is non-convex. We therefore approximate it using convex functions $f:\mathbb{R}\to\mathbb{R}$:
$\mathcal{L}_{f}(\hat{\ell}):=\mathbb{E}\left[f(\hat{\ell}(z)\cdot l)\right].$
In the context of reward modeling, given response pairs $(y_{w},y_{l})$ with preference $y_{w}\succ y_{l}$, we construct binary classification samples with $\ell=1$. For a pointwise reward model $r_{\phi}:\mathcal{Y}\to\mathbb{R}$, the natural parameterization is: ^$\hat{\ell}(y_{w},y_{l})=r_{\phi}(y_{w})-r_{\phi}(y_{l}),$ leading to the loss:
$\mathcal{L}_{f}(\hat{\ell}):=\mathbb{E}\left[f(r_{\phi}(y_{w})-r_{\phi}(y_{l}))\right].$
Different choices of $f(x)$ correspond to distinct loss functions. Specifically, when $f(x)=\mathbb{I}(x<0)$, the 0-1 loss is obtained; when $f(x)=-\log\sigma(x)$, we obtain the logistic loss used in SimPO; and when $f(x)=\text{ReLU}(-x)$, it corresponds to the loss function used in RePO.
@@ 4 Theoretical Analysis of RePO
Let $D\subseteq\mathbb{R}$ be a closed convex set. The convex envelope of $\mathcal{L}_{0\text{-}1}$ over $D$ is:
$\mathrm{conv}_{D}\mathcal{L}_{0\text{-}1}(x)$
::=:= $\sup\left\{h(x)\mid h\text{ is convex},\,h\leq\mathcal{L}_{0\text{-}1}\,\forall x\in D\right\}.$ (17)
Theorem 4.2 (Convex Envelope of 0-1 Loss).
For $D=[-a,b]$ with $a,b>0$, the convex envelope of $\mathcal{L}_{0\text{-}1}(x)=\mathbb{I}(x<0)$ is:
$\mathrm{conv}_{D}\mathcal{L}_{0\text{-}1}(x)=\frac{1}{a}\mathrm{ReLU}(-x).$
@@
  1. 1.
    For binary classification with loss $\mathcal{L}_{0\text{-}1}$:
    $\arg\min_{\hat{\ell}}\mathcal{L}_{0\text{-}1}(\hat{\ell})=\arg\min_{\hat{\ell}}\mathrm{conv}_{D}\mathcal{L}_{0\text{-}1}(\hat{\ell}).$
  2. 2.
    When $D=[-a,b]$:
    $\arg\min_{x\in D}\mathcal{L}_{0\text{-}1}(x)=\arg\min_{x\in D}\frac{1}{a}\mathrm{ReLU}(-x).$
Remark 4.4.
Corollary 4.3 establishes: (i) equivalence between minimizing the 0-1 loss and its convex envelope, (ii) exact optimality preservation for the ReLU surrogate. This guarantees gradient-based optimization converges to solutions matching the theoretical optimum of the intractable 0-1 loss.
Corollary 4.5 (Logistic Loss is Not the Convex Envelope of 0-1 Loss).
@@
Remark 4.6.
Unlike SimPO’s logistic loss (Corollary 4.5), the ReLU-based surrogate inherits the 0-1 loss’s global optimality. This explains RePO’s robustness to suboptimal local minima during gradient-based training.
@@ 5 Experiments
In this section, we present our main experimental findings, demonstrating that RePO consistently achieves better results than existing preference optimization methods across multiple benchmarks.
@@ 5.1 Experimental Setup
The main experimental setup has been described in Section 3.2. We increase a model: Mistral2-7B (Jiang et al., 2023). Additionally, the v0.2 Llama3-Instruct setup uses RLHFlow/ArmoRM-Llama3-8B-v0.1 (Wang et al., 2024b) as the reward model for ranking generated data, significantly enhancing performance. As for baselines, we compare RePO with several state-of-the-art preference optimization methods: DPO (Rafailov et al., 2023), SimPO (Meng et al., 2024), IPO (Azar et al., 2023), CPO (Xu et al., 2024), KTO (Ethayarajh et al., 2024), ORPO (Hong et al., 2024), and R-DPO (Park et al., 2024). We also include the SFT model as a baseline. We thoroughly tune the hyperparameters for each baseline and report the best performance. Further details can be found in Appendix C.1. We also evaluate on downstream tasks from the Huggingface Open Leaderboard benchmarks (Beeching et al., 2023), with additional details in in Appendix C.2. The code is available at https://github.com/junkangwu/RePO.
@@
w. RePO $50.3^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+4.4\%}$ $51.8^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+9.1\%}$ $38.2^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+8.5\%}$ $73.8^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+4.8\%}$ $71.0^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+6.1\%}$ $64.2^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+9.2\%}$
w. RePO++ $50.8^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+5.4\%}$ $52.2^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+9.9\%}$ $37.2^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+5.7\%}$ $71.8^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+2.0\%}$ $69.5^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+3.9\%}$ $65.7^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+11.7\%}$
@@
w. RePO $57.7^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+3.8\%}$ $51.1^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+3.0\%}$ $35.2^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+4.8\%}$ $73.6^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+1.7\%}$ $66.6^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+2.5\%}$ $59.1^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+2.2\%}$
w. RePO++ $56.1^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+0.9\%}$ $50.1^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+1.0\%}$ $35.9^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+6.8\%}$ $74.1^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+2.3\%}$ $66.5^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+2.3\%}$ $59.8^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+3.5\%}$
5.2 Main Results
RePO consistently and significantly outperforms existing preference optimization methods. As shown in Table 1, all preference optimization methods outperform the SFT model. Nevertheless, RePO — despite its simplicity — achieves the strongest overall performance across all benchmarks and settings. The consistent gains highlight its robustness and efficiency. Specifically, RePO surpasses the best-performing baseline by 0.2 to 2.8 points on the AlpacaEval 2 LC win rate under various configurations. On Arena-Hard, RePO generally performs better than competing methods, although CPO achieves slightly higher scores at times. We observe that CPO tends to produce responses that are approximately 50% longer. Because Arena-Hard does not penalize response length, longer outputs may occasionally yield higher evaluations.
5.3 Adaptations of RePO
RePO is a versatile approach that can be adapted to different optimization settings. In this work, we explore two adaptations: RePO and RePO++, applied to both DPO and SimPO. The detailed formulations are provided in Section 3.5.
As shown in Table 2, integrating RePO with DPO and SimPO consistently enhances performance. Notably, RePO++ achieves superior results in most scenarios, which we attribute to its combination of ReLU activation and the original weighting function $s_{\theta}$. This design effectively mitigates overoptimization while preserving the benefits of the original scheme. Interestingly, applying RePO and RePO++ to DPO yields particularly significant improvements (5%–12%), achieving a high score of 65.7 on the Arena-Hard benchmark.
5.4 Dynamic Margin Scheduling
To further investigate the impact of the target reward margin $\gamma$ , we conduct an ablation study using dynamic values for $\gamma$ Specifically, we implement two strategies: (i) $\gamma$ increases from small to large, and (ii) $\gamma$ decreases from large to small. Using RePO, we control the initial and final values of $\gamma$ , adjusting it uniformly over the course of one training epoch. As shown in Figure 7, both excessively small and large values of $\gamma$ lead to suboptimal performance (e.g., , $1.0\rightarrow 0.8$ , $0.0\rightarrow 0.2$ ). In contrast, starting with a moderately large value of $\gamma$ and gradually decreasing it during training (e.g., , $1.0\rightarrow 0.2$ ) results in improved model performance. We interpret this behavior as follows: in the early stages of training, the model is underfitting and thus benefits from stronger updates. In later stages, reducing $\gamma$ helps mitigate overfitting, effectively addressing this issue.
Figure 7: Performance of RePO with dynamic $\gamma$. The dashed line represents RePO with a fixed $\gamma$.

### v2

@@ 3.3 Over-Optimization Analysis
The study of over-optimization can be traced back to traditional RLHF literature, and has been empirically investigated in both controlled experiments [9] and user studies [10]. In this work, we follow their experimental setup to further explore this phenomenon.
Model Over-Optimization: Building on Rafailov et al. [16], we investigate over-optimization in RePO, by evaluating six different values of $\gamma$ ($0.0,0.2,0.4,0.6,0.8,1.0$), each corresponding to varying levels of data filtering. Across all cases, we observe a distinct hump-shaped performance pattern: while moderate filtering improves alignment, excessive filtering causes performance to degrade, highlighting the over-optimization effect.
Figure 6: Over-optimization patterns for RePO on Llama3-8B-IT and Gemma2-9B-IT, using AlpacaEval2 LC win rates and Arena-Hard raw win rates. Dotted curves represent theoretical fits based on Gao et al. [9]’s scaling laws, using GPT-4 win rates instead of standard reward model scores.
Scaling Law Fits. Previous work [9, 16] has established scaling laws for reward model scores as a function of the KL divergence between the initial and optimized policies. In contrast, we eliminate the reference model and the associated computational cost of calculating KL divergence. Instead, we use the mean implicit reward margin during training as a proxy metric. The reward function $R(d)$ is given by:
@@
3.4 RePO++: Exploring Extensions of ReLU-based Filtering
While exploring ReLU’s thresholding behavior, we observed an interesting limitation: for cases where the implicit reward margin is smaller than $\gamma$, their gradient weights become uniform, not differentiating between samples of varying difficulty.
This observation naturally led us to wonder: could we preserve the effective filtering mechanism while reintroducing some degree of weighting? To explore this question, we experimented with combining ReLU’s binary filtering with SimPO’s continuous weighting:
@@ 3.4 RePO++: Exploring Extensions of ReLU-based Filtering
This exploration was a natural follow-up to our main discovery about ReLU’s effectiveness, rather than our primary contribution. We were curious to see whether combining the best aspects of both approaches might yield additional insights about preference learning mechanisms.
What does this combined approach reveal? To understand the behavior of this extension, we examined its gradient with respect to the parameters $\theta$:
@@
where $s_{\theta}=\sigma\left(\beta\left(-M_{\theta}+\gamma\right)\right)$ and $\mathbb{I}_{\theta}$ is an indicator function that is 1 if $M_{\theta}<\gamma$ and 0 otherwise.
We observed that this gradient combines properties we discovered in both approaches: it scales updates by $s_{\theta}$ (similar to SimPO) and filters them using $\mathbb{I}_{\theta}$ (the key discovery in our ReLU exploration), focusing the model on less-separated pairs while giving higher weights to smaller separations.
@@
4 Theoretical Analysis: ReLU’s Optimality in Preference Learning
Next, we establish a surprising theoretical connection between preference optimization and binary classification, revealing why our simple ReLU-based approach achieves superior performance.
Following Tang et al. [17], preference learning can be reformulated as binary classification. Given pairs $(z,l)$ where $z\in\mathbb{R}^{k}$ and $l\in\{-1,1\}$, we aim to learn a predictor $\hat{\ell}(z)$ whose sign matches $l$. The classification accuracy is: $\frac{1}{2}\mathbb{E}[\mathrm{sign}(\hat{\ell}(z)\cdot l)]+\frac{1}{2}.$ This corresponds to minimizing the 0-1 loss:
$\mathcal{L}_{0\text{-}1}(\hat{\ell}):=\mathbb{E}\left[1-\mathrm{sign}(\hat{\ell}(z)\cdot l)\right]$ (14)
For preference data $(y_{w},y_{l})$ where $y_{w}\succ y_{l}$, we set $l=1$ and parameterize $\hat{\ell}(y_{w},y_{l})=r_{\phi}(y_{w})-r_{\phi}(y_{l})$, yielding the objective:
$\mathcal{L}_{f}(\hat{\ell}):=\mathbb{E}\left[f(r_{\phi}(y_{w})-r_{\phi}(y_{l}))\right]$ (15)
Where $f$ determines the surrogate loss: $f(x)=\mathbb{I}(x<0)$ gives the 0-1 loss, $f(x)=-\log\sigma(x)$ yields SimPO’s logistic loss, and $f(x)=\text{ReLU}(-x)$ gives our method’s loss.
Our key insight comes from analyzing the convex envelope of the 0-1 loss:
@@ 4 Theoretical Analysis: ReLU’s Optimality in Preference Learning
The convex envelope of $\mathcal{L}_{0\text{-}1}$ over a closed convex set $D\subseteq\mathbb{R}$ is:
$\mathrm{conv}_{D}\mathcal{L}_{0\text{-}1}(x):=\sup\left\{h(x)\mid h\text{ is convex},\,h\leq\mathcal{L}_{0\text{-}1}\,\forall x\in D\right\}$ (16)
Theorem 4.2 (ReLU as Convex Envelope).
For D=[a,b]D=[-a,b] with $a,b>0$, the convex envelope of $\mathcal{L}_{0\text{-}1}(x)=\mathbb{I}(x<0)$ is:
$\mathrm{conv}_{D}\mathcal{L}_{0\text{-}1}(x)=\frac{1}{a}\mathrm{ReLU}(-x)$ (17)
This remarkable result reveals that ReLU provides the tightest possible convex approximation to the ideal 0-1 loss, explaining its empirical effectiveness. Furthermore:
@@
[…]
And for D=[a,b]D=[-a,b]:
$\arg\min_{x\in D}\mathcal{L}_{0\text{-}1}(x)=\arg\min_{x\in D}\frac{1}{a}\mathrm{ReLU}(-x)$ (19)
This guarantees that gradient-based optimization of our ReLU surrogate converges to solutions matching the theoretical optimum of the intractable 0-1 loss. Importantly:
Corollary 4.4 (Logistic Loss Suboptimality).
@@
This theoretical foundation explains why our simple ReLU-based approach consistently outperforms more complex mechanisms like SimPO’s sigmoid weighting — ReLU provides optimality guarantees that logistic loss cannot match, while being computationally more efficient.
@@ 5 Experiments
In this section, we examine how our simplified ReLU-based approach behaves across different models and settings. Rather than focusing solely on performance gains, we explore patterns that help explain why such a simple mechanism works effectively in practice.
@@ 5.1 Experimental Setup
The core experimental configuration extends our investigation from Section 3.2 to include Mistral2-7B [18] alongside previously examined models. For the Llama3-Instruct v0.2 experiments, we employed the RLHFlow/ArmoRM-Llama3-8B-v0.1 [19] reward model for ranking generated data. We benchmark our approach against established preference optimization methods: DPO [6], SimPO [11], IPO [20], CPO [21], KTO [22], ORPO [23], and R-DPO [24], with SFT models serving as baselines. Implementation details are provided in Appendix D.1. We also evaluate on downstream tasks from the Huggingface Open Leaderboard benchmarks [25], with additional details in in Appendix D.2. The code is available at https://github.com/junkangwu/RePO.
@@
5.2 Result Comparisons
Observation: Simple ReLU thresholding exhibits surprising effectiveness. Table 1 reveals an unexpected pattern: despite removing components previously thought essential, the simple ReLU-based approach consistently performs well across all evaluated models and benchmarks. This finding aligns with our theoretical analysis showing that binary thresholding directly approximates the convex envelope of the 0-1 loss. On AlpacaEval 2, we observe improvements of 0.2-2.8 points in LC win rates across different configurations compared to the strongest baselines.
5.3 Methodology Comparisons
Beyond alignment, we also compare the methodologies of these preference learning methods. Our method plays a hub to connect these methods.
Relation to SimPO. SimPO employs sigmoid weighting via log-sigmoid activation to attenuate gradients as reward margins increase, mitigating over-optimization. RePO can be viewed as “SimPO without log-sigmoid,” replacing this continuous scaling with binary filtering. To validate this relationship, we integrated a ReLU-based filtering mechanism into SimPO (cf. RePO++ Equation 12). Table 2 confirms that ReLU’s filtering mechanism enhances performance. As demonstrated in Section 3.4, RePO++ directly addresses over-optimization while retaining the benefits of some weighting.
Relation to SLiC-HF. RePO can be characterized as “SLiC-HF without SFT regularization”. To ensure a fair comparison (while disregarding differences in length normalization), we investigated the impact of SFT regularization by varying its coefficient, $\lambda$. The results, presented in Appendix Table 6 and further details in Appendix D.5, indicate that this additional regularization term offers no discernible improvement. This suggests that SFT regularization targets a different optimization challenge, distinct from the direct over-optimization problem RePO addresses.
Relation to DPO. Mathematically, DPO is equivalent to SimPO when the margin $\gamma$ is defined as $\log\pi_{\text{ref}}(y_{w}\mid x)-\log\pi_{\text{ref}}(y_{l}\mid x)$ (ignoring length normalization). However, directly substituting the log-sigmoid function with ReLU in DPO’s formulation leads to a significant performance degradation (see Appendix Table 7 and Appendix D.6). This underscores the critical role of the threshold $\gamma$ in determining the effectiveness of over-optimization prevention. As identified by Wu et al. [26], reference model based reward margins are often unreliable as target margins, which explains why SimPO’s explicit $\gamma$ parameter is effective for preference learning.
5.4 Effect of ReLU Filtering Across Methods
Having observed the effectiveness of binary thresholding, we naturally questioned whether this mechanism might enhance other preference learning approaches. Table 2 shows that integrating ReLU filtering consistently improved performance across both DPO and SimPO frameworks, suggesting that selective gradient application based on margin thresholds provides benefits beyond our specific implementation. Our experiments with the combined approach (cf. RePO++ in Section 3.4) revealed particularly strong improvements when applied to DPO (5%–12% gains), with notable performance on Arena-Hard (reaching 65.7). This design effectively mitigates over-optimization while preserving the benefits of the original scheme.
@@ 5.4 Effect of ReLU Filtering Across Methods
w. RePO $50.3^{\color[rgb]{0,0,0}+4.4\%}$ $51.8^{\color[rgb]{0,0,0}+9.1\%}$ $38.2^{\color[rgb]{0,0,0}+8.5\%}$ $73.8^{\color[rgb]{0,0,0}+4.8\%}$ $71.0^{\color[rgb]{0,0,0}+6.1\%}$ $64.2^{\color[rgb]{0,0,0}+9.2\%}$
w. RePO++ $50.8^{\color[rgb]{0,0,0}+5.4\%}$ $52.2^{\color[rgb]{0,0,0}+9.9\%}$ $37.2^{\color[rgb]{0,0,0}+5.7\%}$ $71.8^{\color[rgb]{0,0,0}+2.0\%}$ $69.5^{\color[rgb]{0,0,0}+3.9\%}$ $65.7^{\color[rgb]{0,0,0}+11.7\%}$
@@
w. RePO $57.7^{\color[rgb]{0,0,0}+3.8\%}$ $51.1^{\color[rgb]{0,0,0}+3.0\%}$ $35.2^{\color[rgb]{0,0,0}+4.8\%}$ $73.6^{\color[rgb]{0,0,0}+1.7\%}$ $66.6^{\color[rgb]{0,0,0}+2.5\%}$ $59.1^{\color[rgb]{0,0,0}+2.2\%}$
w. RePO++ $56.1^{\color[rgb]{0,0,0}+0.9\%}$ $50.1^{\color[rgb]{0,0,0}+1.0\%}$ $35.9^{\color[rgb]{0,0,0}+6.8\%}$ $74.1^{\color[rgb]{0,0,0}+2.3\%}$ $66.5^{\color[rgb]{0,0,0}+2.3\%}$ $59.8^{\color[rgb]{0,0,0}+3.5\%}$
5.5 Dynamic Margin Scheduling and Curriculum Learning
Our investigation into the role of target reward margin $\gamma$ led us to an unexpected discovery about curriculum learning. We experimented with dynamic scheduling of $\gamma$ throughout training, implementing two strategies: (i) increasing $\gamma$ from small to large, and (ii) decreasing $\gamma$ from large to small. Figure 7 reveals a striking pattern: starting with a moderately large value of $\gamma$ and gradually decreasing it ($1.0\rightarrow 0.2$) naturally creates an effective curriculum that improves model performance. In contrast, both excessively large values ($1.0\rightarrow 0.8$) and small values ($0.0\rightarrow 0.2$) led to suboptimal outcomes.
Figure 7: Exploration of dynamic $\gamma$ scheduling on Llama3-Instruct v0.2 (8B). The dashed line represents performance with a fixed $\gamma$. We observed that decreasing $\gamma$ from an initially larger value creates a natural curriculum that enhances performance.
This observation reveals an intriguing self-regulating property: early in training when the model is underfitting, a larger $\gamma$ permits more aggressive updates across more examples. As training progresses, the decreasing $\gamma$ naturally focuses learning on increasingly challenging examples, effectively preventing over-optimization. This emergent curriculum behavior, arising from a simple parameter schedule, suggests that binary thresholding captures fundamental learning dynamics that more complex mechanisms might obscure.